Compute CPl of the following mix of instructions. Instruction type Instruction-mix in a program of 100 instructions 15 25 15 35 Cycles Store Load Branch Integer arithmetic Integer shift 1 2 4 1 1 10 Integer multi
Answer #1

CPI = sigma{i}(IC{i}*CC{i} / IC

Where {\displaystyle IC_{i}}{\displaystyle IC_{i}} is the number of instructions for a given instruction type {\displaystyle i}i, {\displaystyle CC_{i}}{\displaystyle CC_{i}} is the clock-cycles for that instruction type and {\displaystyle IC=\Sigma _{i}(IC_{i})}{\displaystyle IC=\Sigma _{i}(IC_{i})} is the total instruction count. The summation sums over all instruction types for a given benchmarking process.

CPI = (15*1 + 25*2 + 15*4 + 35*1 + 5*1 + 5*10)/100 = 2.15 ans